OpenStudy (anonymous):

What is the solution to the proportion? 3y-8/12 = y/5

OpenStudy (anonymous):

simplify i.e 3y-2/3=y/5 multiply by 5 thruogh out i.e 15y -10/3=y simplify father y=2/9

OpenStudy (tkhunny):

\(3y - \dfrac{8}{12} = \dfrac{y}{5}\) is NOT a proportion. If you intended to write \(\dfrac{3y-8}{12} = \dfrac{y}{5}\), please remember your Order of Operations and give it another go. Use parentheses where necessary.